OpenStudy (anonymous):

0= cos^2 theta - sin ^2 theta how do i know whether i should switch cos^2 theta to (1-sin^2 theta) or switch sin^2 theta to ( 1- cos^2 theta) both ways work but give a different answer

OpenStudy (anonymous):

im guessing this is for integral of sin/cos -- ( it all depends wat u get, sometime it will cost more work if u used the other one)

OpenStudy (anonymous):

sin^2(theta)=cos^2(theta) implies tan^2(theta)=1 implies tan(theta)= +-1 implies theta = arctan(+-1) implies theta = +- pi/4, which you can see is a solution upon inspection.

OpenStudy (anonymous):

Technically there are infinite answers, but this is two of them.

OpenStudy (callisto):

I was thinking..\[cos^2 \theta - sin ^2 \theta = cos2\theta\]

OpenStudy (anonymous):

i need to solve for theta..i dont think cos2theta will have the answer of 45deg

OpenStudy (anonymous):

but technically, all ways should work out isnt it..i dont understand why it gives different answers ..so how do i know when to use what

OpenStudy (anonymous):

@iHelp but both ways give different answers...so is not more work..or did i do something wrong..

OpenStudy (anonymous):

I did solve for theta...why does no one see that?

OpenStudy (anonymous):

because dont really understand what u said..

OpenStudy (callisto):

Actually... All the above method gives you the answer 45 ...

OpenStudy (anonymous):

And 45 is an answer because cos(45)=sqrt(2)/2=sin(45) implies sin^2(45)-cos^2(45)=0 Which is your equation...

OpenStudy (anonymous):

So I'm right...

OpenStudy (anonymous):

how is sin theta = cos theta... yea.. the answer is 45deg but i want to know why when i do 0 = cos^2 theta - ( 1- cos^2 theta) it doesnt give 45 as the answer

OpenStudy (anonymous):

cuz sin^2 theta + cos^2 theta = 1 sin^2 theta = 1 - cos^2 theta so my equation should work out too

OpenStudy (callisto):

Method I:\[cos^2\theta-sin^2\theta =0\]\[1-sin^2\theta-sin^2\theta =0\]\[1-2sin^2\theta =0\]\[sin^2\theta =1/2\]\[sin\theta =\sqrt{1/2}\]\[\theta =45\]

OpenStudy (callisto):

Method II \[cos^2\theta-sin^2\theta =0\]\[cosn^2\theta-(1-cos^2\theta) =0\]\[2cos^2\theta -1=0\]\[cos^2\theta =1/2\]\[cos\theta =\sqrt{1/2}\]\[\theta =45\]

OpenStudy (anonymous):

ohh...i find the angle using 1/2 first and then divide it by 2, i caught my mistake thanks!

OpenStudy (callisto):

Method III: \[cos^2\theta-sin^2\theta =0\]\[cos2\theta =0\]\[2\theta =90\]\[\theta =45\]

OpenStudy (anonymous):

wow thanks so much. my doubts are clear (:

OpenStudy (callisto):

Yay~ Glad to hear :)

OpenStudy (callisto):

As you see, method III is the fastest :P just kidding :)